you are placing 12 pictures on separae pages in an album. how many different ways can you order the 12 pictures in the album? how many different ways can 4 of the 12 pictures be placed on the first 4 pages?
Answer by stanbon(75887)

you are placing 12 pictures on separate pages in an album. how many different ways can you order the 12 pictures in the album?
Ans: 12! = 479,001,600
=============================
how many different ways can 4 of the 12 pictures be placed on the first 4 pages?
------------------------------------
12P4 = 12!/(12-4)! = 12*11*10*9 = 11,880
=========================================
Cheers,
Stan H.